If a 0.3 kg object hanging from a spring stretches it by 0.25 m, then by how much will the spring be stretched (in m) if a 0.6 kg object is suspended from it?

m
Answer #1

as spring constant is given by,

k = F / x

where k is spring constant, F is force applied and x is the displacement of spring system.

now for 0.3 kg mass spring stretches 0.25 m.so,

k = ( m * g)/x

k = ( 0.3 * 9.81)/0.25

k = 11.772 N/m

now when the mass is replaced from 0.3 kg to 0.6 kg,

x = (m * g)/k

x = (0.6 * 9.81)/ 11.772

x = 0.5 m

so, the spring will stretch to 0.5 m if a 0.6 kg object is suspended from it.

→ if you find this answer helpful please rate positive thank you so much.